ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: X Rayes, Mohamed Omar, Texas Instruments Incorporated, Dallas, Texas; Rosado, Fernando, Universidade de Lisboa, Portugal On the Outlier Problem for Structured Data The general examination of outliers in a sample of observations suppose a specific distribution. It is assumed the null hypothesis of homogeneity of the distribution. In many statistical analyses, departure from the asumption of homogeneity of distribution need have nothing to do with outliers - and this is the natural manifestationof the appropriateness of some linear models. For ore structured data the concept of outlier is more nebulous. We present a natural war of “how to define” outlier in one general discordancy model.